Provide reagents to achieve the shown multi-step transformations. (6 pts) H Ph Ph Ph. Me
Answer #1

Ph Br Brz/ecla ph- CH -cth 2 equivalent 7 Br NaNt NaNt ph-c=EN Ph-c=c- CEC-H Brt CH2 -CH=en-me ph-ese-CH-CH = CH-Me Lux 42/Li